# Normal coactions extend to the C*-envelope

### 13 Mar 2024 v2

@@ Abstract.
We show that a normal coaction of a discrete group on an operator algebra extends to a normal coaction on the C*-envelope. This resolves an open problem considered by Kakariadis, Katsoulis, Laca, and X. Li, and provides a direct proof of a prominent result of Sehnem. As applications, we resolve a question of X. Li, where we identify the C*-envelopes of the operator algebras of groupoid-embeddable categories and of cancellative right LCM monoids. This latter class includes many examples of monoids that are not group-embeddable.
@@
K.A. Brix was supported by a DFF-international postdoc grant 1025-00004B. C. Bruce has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101022531. A. Dor-On was partially supported by EU grant HORIZON-MSCA-SE-2021 Project 101086394 and by Banff International Research Station for the 2020 Focused Research Group 248. This research was partially supported by a European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme 817597 and the London Mathematical Society through a Research in Pairs Grant 42204.
@@ 1. Introduction
We break new terrain in non-commutative boundary theory by proving that every normal coaction of a discrete group on an operator algebra extends to a coaction on the C*-envelope (Theorem A). Our proof relies only on elementary principles and marks the first use of non-abelian duality theory for coactions on operator algebras, which is inspired by the theory of non-abelian duality for coactions on C*-algebras, as developed by Echterhoff, Kaliszewski, Quigg, and Raeburn, to name a few. This resolves an open problem considered by Kakariadis, Katsoulis, Laca, and X. Li, and yields a short, conceptual proof of a prominent result of Sehnem. As a consequence of our results, the C*-envelope always coincides with the equivariant C*-envelope for a normal coaction (Corollary 3.6), thereby rendering the latter superfluous for normal coactions.
@@
The C*-envelope $C_{\textup{env}}^{*}(\mathcal{A})$ of an operator algebra $\mathcal{A}$ is the smallest C*-algebra containing a completely isometric copy of $\mathcal{A}$ (see § 2.1 for the precise definition). The structure and invariants of the C*-envelope then provide a new lens for studying non-selfadjoint operator algebras via selfadjoint (i.e., C*-algebra) theory. Arveson’s works provide a deep and fruitful connection between the theories of selfadjoint and non-selfadjoint operator algebras, and it is now well-established that structural results for C*-envelopes are crucial for discovering and making use of new invariants for the underlying operator algebra. Indeed, in [DFK17, DK20, DKKLL22, Seh22] non-self-adjoint techniques are used to establish gauge-compatible co-universal properties for C*-algebras, in [KR19, Ka17, DK20, KR21, DKKLL22] the identification of the C*-envelope is used to resolve crossed product Hao–Ng isomorphism problems, and in [DEG20] $K$-theory for C*-algebras is used to classify operator algebras arising from directed graphs up to different notions of isomorphisms.
Beyond interactions with C*-algebra theory, C*-envelopes appear in relation to important open problems in non-commutative boundary theory. These include the Arveson–Douglas essential normality conjecture [Arv05] and Arveson’s hyperrigidity conjecture [Arv11], with applications to the theory of analytic varieties and approximation theory of functions. With such potential applications in mind, Arveson initiated a program of finding explicit descriptions for C*-envelopes of naturally occurring non-selfadjoint operator algebras. A general context in which one can answer this question normally consists of a Toeplitz C*-algebra $\mathcal{T}$ considered as a C*-cover of a canonical norm-closed subalgebra $\mathcal{T}^{+}$ both generated by operators that comprise an analogue of a left regular representation. This non-selfadjoint operator algebra $\mathcal{T}^{+}$ is often called the tensor algebra of the context, and a general scheme for computing its C*-envelope treats several classes of examples of operator algebras simultaneously.
@@
Despite some developments, the lack of Pontryagin duality for non-abelian groups posed a significant challenge for determining the C*-envelope of tensor algebras in Fowler’s context beyond monoids that embed in abelian groups (see [KKLL22a, KKLL22b]). This problem is intimately related to coactions on operator algebras. More precisely, all tensor algebras mentioned above come equipped with a canonical normal coaction of a discrete group (see § 2.1 for the precise definitions). It was observed in [DK20] that the problem of computing the C*-envelope in Fowler’s context can be abstracted to the more general problem of whether the canonical coaction on an operator algebra extends to its C*-envelope.
@@
Progress on this question in Fowler’s context for normal coactions (e.g., all coactions by amenable groups) was made by Kakariadis, Katsoulis, Laca, and X. Li [KKLL22a, KKLL22b] under additional assumptions, and Sehnem resolved the problem for product systems in [Seh22]. More precisely, Sehnem showed that the covariance algebra from [Seh18] coincides with the C*-envelope for product systems over an arbitrary group-embeddable monoid, circumventing Question A for general normal coactions. Sehnem’s paper [Seh22] was a major advance and marks the state-of-the-art, but the proofs require intricate technical arguments in the specific setting of product systems leaving Question A open in general, see [Seh22, Remark 5.2]. Our main contribution is to resolve Question A for all normal coactions.
@@
Let $\mathcal{A}$ be an operator algebra with a contractive approximate identity. Then, any normal coaction of a discrete group $G$ on $\mathcal{A}$ has a (necessarily unique) extension to a normal coaction of $G$ on $C_{\textup{env}}^{*}(\mathcal{A})$.
By the co-universal property of the C*-envelope, every action of a discrete group on an operator algebra extends to an action on the C*-envelope. However, it is not at all clear whether a coaction extends to the C*-envelope. In order to overcome this problem, we extend Katayama duality [Ka84] to normal coactions of discrete groups on operator algebras (Theorem 3.3). This allows us to embed the C*-envelope in a reduced double crossed product in such a way that the restriction of the canonical coaction on the double crossed product gives us the desired coaction. This approach is concise, and leads to new avenues of research. For instance, Theorem A may lead to new applications in dilation theory, and has important consequences for Arveson’s program of computing C*-envelopes.
First, in the special case of product systems over group-embeddable monoids, Theorem A leads to a short, conceptually simpler proof of Sehnem’s main result, [Seh22, Theorem 5.1] (thus also recovering results in [KKLL22a, KKLL22b]). Second, for an operator algebra arising from a cancellative small category, we use Theorem A to provide a sufficient condition for its C*-envelope to coincide with the boundary quotient C*-algebra of the category. We verify this sufficient condition in two important examples classes. In the case when the small category is groupoid-embeddable or in the case when the category is a cancellative (not necessarily group-embeddable) right LCM monoid, which then yields a description of the C*-envelope, see Theorems B and C. We emphasize that examples of operator algebras arising from cancellative small categories do not naturally fit into Fowler’s context. On the other hand, examples of operator algebras that do fit in that context now enjoy a significantly more conceptual proof for computing their C*-envelope, avoiding the technical language of product systems. As this application makes novel use of étale groupoid theory in the setting of non-selfadjoint operator algebras, we next explain the context and consequences of our applications.
@@
Our final application is to identify the C*-envelopes of the operator algebras of all cancellative right LCM monoids. Importantly, this class contains Dehornoy’s gcd-monoids [Doh17I, Doh17I], where there are concrete examples of monoids that are not group-embeddable, see [DW17, Proposition 4.3]. Therefore, this class of monoids is not covered by any previous results, including [Seh22].
@@
Acknowledgements. We are grateful to Xin Li for discussions on non-abelian duality that led to simplified proofs, and for bringing the reference [DW17] to our attention. The third-named author acknowledges the March 2020 FRG 248 BIRS meeting in Banff.
@@ 2.1. Coactions on operator algebras
We first discuss operator algebras and their C*-envelopes. The reader is referred to [Arv69, Arv72, Arv] for the unital theory, and to [DS18, § 2.2] for additional details on the general theory.
By an operator algebra $\mathcal{A}$, we mean a norm-closed subalgebra of the bounded operators $\mathbb{B}(\mathcal{H})$ on a Hilbert space $\mathcal{H}$. A C*-cover of an operator algebra $\mathcal{A}$ is a pair $(\iota,\mathcal{B})$, where $\mathcal{B}$ is a C*-algebra and $\iota\colon\mathcal{A}\to B$ is a completely isometric homomorphism such that $\mathcal{B}=C^{*}(\iota(\mathcal{A}))$. The C*-envelope of $\mathcal{A}$ is a C*-cover $(\kappa,C^{*}_{\textup{env}}(\mathcal{A}))$ of $\mathcal{A}$ with the co-universal property that for every other C*-cover $(\iota,\mathcal{B})$ of $\mathcal{A}$, there is a surjective *-homomorphism $q_{e}\colon C^{*}(\mathcal{B})\rightarrow C^{*}_{\textup{env}}(\mathcal{A})$ with $q_{e}\circ\iota=\kappa$ on $\mathcal{A}$. By this co-universal property, the C*-algebra $C^{*}_{\textup{env}}(\mathcal{A})$ is unique up to canonical *-isomorphism.
@@
$\|a\otimes I-\sum_{g}\delta(a_{g})(I_{\mathcal{H}}\otimes u_{g^{-1}})\|<\epsilon.$
Since $\textup{id}_{\mathcal{A}}\otimes\Phi_{e}$ is contractive, we get in $\mathcal{A}\otimes\mathbb{C}u_{e}$ that
$\|a\otimes I-\sum_{g}\mathbb{E}(a_{g})(I_{\mathcal{H}}\otimes u_{g^{-1}})\|<\epsilon.$
By the coaction identity, we have $\mathbb{E}(a_{g})(I_{\mathcal{H}}\otimes u_{g^{-1}})\in\mathcal{A}_{g}\otimes\mathbb{C}I$, so $a$ can be approximated up to arbitrary tolerance by an element in $\sum_{g}\mathcal{A}_{g}$. ∎
@@
It follows from Proposition 2.3 that any coaction $\delta$ is automatically completely isometric. Indeed, if we let $1\colon G\rightarrow\mathbb{C}$ be the trivial representation, then we have $(\textup{id}_{\mathcal{A}}\otimes 1)\circ\delta=\textup{id}_{\mathcal{A}}$, as can be verified on on spectral subspaces.
@@
For the proof of Theorem 3.5, we need a slight strengthening of [DKKLL22, Proposition 3.4] which uses Fell’s absorption principle. See [CD23, § 5.5] and [Ka23] for versions of Fell’s absorption principle for monoid operator algebras.
@@ 2.2. Operator algebras from categories
The operator algebra $\mathcal{A}_{\lambda}(\mathfrak{C})$ is unital whenever $\mathfrak{C}^{0}$ is finite and has a contractive approximate identity comprised of projections. Indeed, the net of projections $\sum_{u\in F}\lambda_{u}$ as $F$ ranges over the finite subsets of $\mathfrak{C}^{0}$ is a contractive approximate identity of projections for the dense algebra $\textup{alg}(\{\lambda_{\mathfrak{C}}(c):c\in\mathfrak{C}\})$, so it is also a contractive approximate identity for $\mathcal{A}_{\lambda}(\mathfrak{C})$.
@@
where the product $c_{1}^{-1}d_{1}\cdots c_{n}^{-1}d_{n}$ is taken inside $\mathcal{I}(\mathfrak{C})$. We denote by 00 the empty function on $\mathfrak{C}$. Using [CELY17, Lemma 5.6.43], we see that $I_{l}$ contains 00 if and only if $\mathfrak{C}$ is a monoid such that $x\mathfrak{C}\cap y\mathfrak{C}\neq\emptyset$ for all $x,y\in\mathfrak{C}$. Combining (2.1) and (2.4), we get that
@@
if there exists $X\in\mathcal{J}$ such that $\chi(X)=1$ and $s(x)=t(x)$ for all $x\in X$. We let $[s,\chi]$ denote the equivalence class of $(s,\chi)$ with respect to $\sim$. The transformation groupoid $I_{l}\ltimes\Omega$ is the quotient space $(I_{l}*\Omega)/\sim$ with groupoid operations determined by
$[s,t.\chi][t,\chi]\coloneq[st,\chi]\quad\text{ and }\quad[s,\chi]^{-1}\coloneq[s^{-1},s.\chi]$
for all $s,t\in I_{l}$ and $\chi\in\Omega$. The range and source maps are given by $\mathfrak{r}([s,\chi])\coloneq s.\chi$ and $\mathfrak{s}([s,\chi])\coloneq\chi$, respectively, for all $[s,\chi]\in I_{l}\ltimes\Omega$. We tacitly identify $\Omega$ with the unit space of $I_{l}\ltimes\Omega$ via $\chi\mapsto[\textup{id}_{X},\chi]$ for all $\chi\in\Omega$, where $X\in\mathcal{J}$ is any constructible right ideal with $\chi(X)=1$. Subsets of the form $[s,U]\coloneq\{[s,\chi]:\chi\in U\}$ for all $s\in I_{l}$ and compact open subsets $U$ of $\Omega(\textup{dom}(s))$ generate a topology on the groupoid $I_{l}\ltimes\Omega$ which is locally compact, étale (both range and source maps are local homeomorphisms), and ample (there is a basis consisting of compact open bisections) though not necessarily Hausdorff. A bisection is a subset of $I_{l}\ltimes\Omega$ on which both the range and source maps are injective, and any basic open subset $[s,U]$ as above is a compact open bisection.
@@
Let us recall some terminology from [Ex08, Definition 11.5]. A subset $F\subseteq\mathcal{J}$ of constructible right ideals is said to be a cover for $X\in\mathcal{J}$ if $Z\subseteq X$ for all $Z\in F$ and for every $Y\in\mathcal{J}$ with $Y\subseteq X$, there exists $Z\in F$ such that $Z\cap Y\neq\emptyset$. If $0\in I_{l}$, then $\partial\Omega$ is precisely the set of tight characters of $\mathcal{J}$ in the sense of Exel, see [Ex08, Theorem 12.9]. Precisely, this means that $\chi\in\Omega$ lies in $\partial\Omega$ if and only if whenever $X\in\mathcal{J}^{\times}$ and $F$ is a finite cover for $X$, we have $\chi(Y)=1$ for some $Y\in F$. This implies that $\partial\Omega(X,\mathfrak{f})=\emptyset$ whenever $\mathfrak{f}$ is a cover for $X$.
@@
for all $f\in C_{c}(I_{l}\ltimes\Omega)$ and $[t,\chi]\in(I_{l}\ltimes\Omega)_{\chi}$. Similarly, given $\chi\in\partial\Omega$, we let
@@
determined by $q_{\partial}(f)=f|_{I_{l}\ltimes\partial\Omega}$ for all $f\in C_{c}(I_{l}\ltimes\Omega)$.
When $I_{l}\ltimes\Omega$ is Hausdorff, the isomorphism in (2.6) justifies Definition 2.11. In the non-Hausdorff setting, there is another candidate for a groupoid model for $C_{\lambda}^{*}(\mathfrak{C})$ (see [Li23, § 3]), and we do not know if the two groupoids are different (see [Li23, Question 3.6]).

### 17 Mar 2026 v3

@@ Abstract.
We show that a normal coaction of a discrete group on an operator algebra extends to a normal coaction on the C*-envelope. This resolves an open problem attempted by several experts in the area, and provides a more direct proof of a prominent result of Sehnem. As an application, we resolve a question of X. Li, where we identify the C*-envelopes of the operator algebras of groupoid-embeddable categories and of cancellative right LCM monoids. This latter class includes many examples of monoids that are not group-embeddable.
@@
K.A. Brix was supported by a DFF-international postdoc grant 1025-00004B. C. Bruce has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101022531. A. Dor-On was partially supported by an NSF / BSF grant no. 2530543 / 2023695 (respectively) and a DFG Middle-Eastern collaboration project no. 529300231. This research was partially supported by a European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme 817597 and the London Mathematical Society through a Research in Pairs Grant 42204.
@@ 1. Introduction
In this paper, we show that every normal coaction of a discrete group on an operator algebra extends to a coaction on the C*-envelope (Theorem A). Our proof relies only on elementary principles from non-abelian duality theory for coactions on operator algebras inspired by the theory of non-abelian duality for coactions on C*-algebras as developed by Echterhoff, Kaliszewski, Quigg, and Raeburn [EKQR06], as well as duality theory for actions by abelian groups on operator algebras as developed by Katsoulis and Ramsey [KR19, KR18], to name but a few. Our work resolves an open problem considered by several experts in the area, and yields a shorter, conceptual proof of a prominent result of Sehnem. As a consequence of our results, the C*-envelope always coincides with the equivariant C*-envelope for a normal coaction (Corollary 3.6).
@@
The C*-envelope $C_{\textup{env}}^{*}(\mathcal{A})$ of an operator algebra $\mathcal{A}$ is the smallest C*-algebra containing a completely isometric copy of $\mathcal{A}$ (see § 2.1 for the precise definition). The structure and invariants of the C*-envelope then provide a new lens for studying non-selfadjoint operator algebras via selfadjoint (i.e., C*-algebra) theory. Arveson’s works provide a deep and fruitful connection between the theories of selfadjoint and non-selfadjoint operator algebras, and it is now well-established that structural results for C*-envelopes are crucial for discovering and making use of new invariants for the underlying operator algebra. Indeed, in [DFK17, DK20, DKKLL22, Seh22] non-self-adjoint techniques are used to establish gauge-compatible co-universal properties for C*-algebras, in [KR19, Ka17, DK20, KR21] the identification of the C*-envelope (or the coaction C*-envelope in [DKKLL22]) is used to resolve crossed product Hao–Ng isomorphism problems, and in [DEG20] the computation of the C*-envelope together with $K$-theory for C*-algebras are used to classify operator algebras arising from directed graphs up to different notions of isomorphisms.
Beyond interactions with C*-algebra theory, C*-envelopes appear in relation to important open problems in non-commutative boundary theory. These include the Arveson–Douglas essential normality conjecture [Arv05] and Arveson’s hyperrigidity conjecture [Arv11] (which was recently resolved by the third-named author together with Bilich in [BD+]), with applications to the theory of analytic varieties and approximation theory of functions. With such potential applications in mind, Arveson initiated a program of finding explicit descriptions for C*-envelopes of naturally occurring non-selfadjoint operator algebras. A general context in which one can answer this question normally consists of a Toeplitz C*-algebra $\mathcal{T}$ considered as a C*-cover of a canonical norm-closed subalgebra $\mathcal{T}^{+}$ both generated by operators that comprise an analogue of a left regular representation. This non-selfadjoint operator algebra $\mathcal{T}^{+}$ is often called the tensor algebra of the context, and a general scheme for computing its C*-envelope treats several classes of operator algebras simultaneously.
@@
Despite some developments, the lack of Pontryagin duality for non-abelian groups posed a significant challenge for determining the C*-envelope of tensor algebras in Fowler’s context beyond monoids that embed in abelian groups (see [DKKLL22, KKLL22a, KKLL22b]). This problem is intimately related to coactions on operator algebras. More precisely, all tensor algebras mentioned above come equipped with a canonical normal coaction of a discrete group (see § 2.1 for the precise definitions). It was observed in [DK20] that the problem of computing the C*-envelope in Fowler’s context can be abstracted to the more general problem of whether the canonical coaction on an operator algebra extends to its C*-envelope.
@@
Progress on this question in Fowler’s context for normal coactions (e.g., all coactions by amenable groups) was made in [KKLL22a, KKLL22b] under additional assumptions, and was eventually resolved completely for group-embeddable monoids by Sehnem in [Seh22]. More precisely, Sehnem showed that the covariance algebra from [Seh18] coincides with the C*-envelope for product systems over an arbitrary group-embeddable monoid, circumventing Question A for general normal coactions. Sehnem’s paper [Seh22] was a major advance and marks the state-of-the-art, but the proofs require intricate technical arguments in the specific setting of product systems, leaving Question A open in general (see [Seh22, Remark 5.2]).
Duality theory for operator algebras has a long history in the context of C*-algebras, dating back to the work of Takesaki and Takai in the 1970s [Tak73, Tak75]. Surprisingly, this theory has found fertile grounds only recently in the general theory of operator algebras with the work of Katsoulis and Ramsey [KR19], where a version of Takai duality for general operator algebras was proven. This led to applications in establishing semisimplicity of certain non-selfadjoint operator algebras. Further applications of duality theory of operator algebras are found in [KR18, EKR25]. Thus, duality theory is emerging as a powerful tool for the study of non-selfadjoint operator algebras. One of our main contributions in this paper is to showcase the strength of non-abelian duality theory of operator algebras to resolve Question A for all normal coactions.
@@
Let $\mathcal{A}$ be an operator algebra with a contractive self-adjoint approximate identity. Then, any normal coaction of a discrete group $G$ on $\mathcal{A}$ has a (necessarily unique) extension to a normal coaction of $G$ on $C_{\textup{env}}^{*}(\mathcal{A})$.
By the co-universal property of the C*-envelope, every action of a discrete group on an operator algebra extends to an action on the C*-envelope. However, it is not at all clear whether a coaction extends to the C*-envelope. In order to overcome this problem, we extend Katayama duality [Ka84] to normal coactions of discrete groups on operator algebras (Theorem 3.3). This allows us to embed the C*-envelope in a reduced double crossed product in such a way that the restriction of the canonical coaction on the double crossed product gives us the desired coaction. We provide an outline of the proof.
Suppose $\delta\colon G\downtouparrow\mathcal{A}$ is a normal coaction of a discrete group $G$ on an operator algebra $\mathcal{A}$. To show that $C_{\textup{env}}^{*}(\mathcal{A})$ carries a normal coaction of $G$ that extends $\delta$, we consider the embedding of $C_{\textup{env}}^{*}(\mathcal{A})$ in $C^{*}_{\textup{env}}(\mathcal{A})\otimes\mathbb{K}$ by tensoring with a rank-one projection on the second tensor factor, where $\mathbb{K}=\mathbb{K}(\ell^{2}(G))$. By Katayama duality for operator algebras (Theorem 3.3), $\mathcal{A}\otimes\mathbb{K}$ is equivariantly isomorphic to a crossed product by $G$. Now by [Ka17, Theorem 2.5], $C_{\textup{env}}^{*}(\mathcal{A}\otimes\mathbb{K})\cong C^{*}_{\textup{env}}(\mathcal{A})\otimes\mathbb{K}$ is also a reduced crossed product by $G$, and thus carries a canonical normal coaction of $G$. We then prove that $C_{\textup{env}}^{*}(\mathcal{A})$ is invariant under this coaction, so that the restriction of this coaction to the image of $C_{\textup{env}}^{*}(\mathcal{A})$ gives us a normal coaction on $C_{\textup{env}}^{*}(\mathcal{A})$. We then verify on generators that this restriction is an extension of $\delta$.
Our approach is concise, and leads to new avenues of research. For instance, Theorem A may lead to new applications in dilation theory, and has important consequences for Arveson’s program of computing C*-envelopes.
First, in the special case of product systems over group-embeddable monoids, we show in Remark 3.7 how Theorem A leads to a shorter, conceptually simpler proof of Sehnem’s main result, [Seh22, Theorem 5.1] (thus also recovering results in [KKLL22a, KKLL22b]). Second, for an operator algebra arising from a cancellative small category, we use Theorem A to provide a sufficient condition for its C*-envelope to coincide with the boundary quotient C*-algebra of the category. We verify this sufficient condition in two important examples classes. In the case when the small category is groupoid-embeddable or in the case when the category is a cancellative (not necessarily group-embeddable) right LCM monoid, which then yields a description of the C*-envelope, see Theorems B and C. We emphasize that examples of operator algebras arising from cancellative small categories do not naturally fit into Fowler’s context. On the other hand, examples of operator algebras that do fit in that context now enjoy a significantly more conceptual proof for computing their C*-envelope, avoiding the technical language of product systems. As this application makes novel use of étale groupoid theory in the setting of non-selfadjoint operator algebras, we explain the context and consequences of our applications below.
@@
Our final application is to identify the C*-envelopes of the operator algebras of all cancellative right LCM monoids. Importantly, this class contains Dehornoy’s gcd-monoids [Doh17I, Doh17I] where there are concrete examples of monoids that are not group-embeddable (see [DW17, Proposition 4.3]), as well as non group-embeddable Malcev type monoids from [EH25]. Therefore, these classes of monoids are not covered by any previous results, including [Seh22].
@@
Acknowledgements. We thank Xin Li for discussions on non-abelian duality that led to simplified proofs, and for bringing the reference [DW17] to our attention. We are also grateful to Cédric Arhancet for pointing out to us a small gap in the proof of Proposition 2.3 in an earlier version of this paper, and we thank the anonymous referees for several helpful suggestions. The third-named author acknowledges discussions held at the March 2020 FRG 248 BIRS meeting in Banff.
@@ 2.1. Coactions on operator algebras
We first discuss operator algebras and their C*-envelopes. The reader is referred to [Arv69, Arv72, Arv08, Arv11] for the unital theory, and to [DS18, § 2.2] for additional details on the general theory.
By an operator algebra $\mathcal{A}$, we mean a norm-closed subalgebra of the bounded operators $\mathbb{B}(\mathcal{H})$ on a Hilbert space $\mathcal{H}$. A C*-cover of an operator algebra $\mathcal{A}$ is a pair $(\iota,\mathcal{B})$, where $\mathcal{B}$ is a C*-algebra and $\iota\colon\mathcal{A}\to\mathcal{B}$ is a completely isometric homomorphism such that $\mathcal{B}=C^{*}(\iota(\mathcal{A}))$. The C*-envelope of $\mathcal{A}$ is a C*-cover $(\kappa,C^{*}_{\textup{env}}(\mathcal{A}))$ of $\mathcal{A}$ with the co-universal property that for every other C*-cover $(\iota,\mathcal{B})$ of $\mathcal{A}$, there is a surjective *-homomorphism $q_{e}\colon C^{*}(\mathcal{B})\rightarrow C^{*}_{\textup{env}}(\mathcal{A})$ with $q_{e}\circ\iota=\kappa$ on $\mathcal{A}$. By this co-universal property, the C*-algebra $C^{*}_{\textup{env}}(\mathcal{A})$ is unique up to canonical *-isomorphism.
@@
(2.1) $\|a\otimes I-\sum_{g}\delta(a_{g})(I_{\mathcal{H}}\otimes u_{g^{-1}})\|<\epsilon.$
The formula $\Phi_{g}(x)=\Phi_{e}(xu_{g^{-1}})$ for all $x\in C^{*}(G)$ promotes to the following: for all $T\in\mathcal{A}\otimes C^{*}(G)$, we have
$(\textup{id}_{\mathcal{A}}\otimes\Phi_{g})(T)=(\textup{id}_{\mathcal{A}}\otimes\Phi_{e})(T(I_{H}\otimes u_{g^{-1}})).$
This is checked first on elementary tensors, which then yields the general formula by continuity. Using this, we have
$(\textup{id}_{A}\otimes\Phi_{e})(\delta(a_{g})(I_{H}\otimes u_{g^{-1}}))=(\textup{id}_{A}\otimes\Phi_{g})(\delta(a_{g}))=\mathbb{E}_{g}(a_{g}).$
Since $\textup{id}_{\mathcal{A}}\otimes\Phi_{e}$ is contractive, using Equation (2.1), we get that
$\|a\otimes I-\sum_{g}\mathbb{E}_{g}(a_{g})\|<\epsilon.$
Since $\mathbb{E}_{g}(a_{g})\in\mathcal{A}_{g}$, we get that $a$ can be approximated arbitrarily well by an element in $\sum_{g}\mathcal{A}_{g}$. ∎
@@
It follows from Proposition 2.3 that any coaction $\delta$ is automatically completely isometric. Indeed, if we let $1\colon G\rightarrow\mathbb{C}$ be the trivial representation, and we continue to write $1\colon C^{*}(G)\rightarrow\mathbb{C}$ for the incuded representation on $C^{*}(G)$, then we have $(\textup{id}_{\mathcal{A}}\otimes 1)\circ\delta=\textup{id}_{\mathcal{A}}$, as can be verified on spectral subspaces.
Let $\mathcal{A}$ be an operator algebra, let $G$ be a discrete group, and suppose $\delta\colon\mathcal{A}\rightarrow\mathcal{A}\otimes C^{*}(G)$ is a normal coaction. Several times throughout this paper we shall use the fact that, in this case, the conditional expectation $\mathbb{E}_{e}\colon\mathcal{A}\to\mathcal{A}$ is faithful. Indeed, we have $\mathbb{E}_{e}=(\textup{id}_{\mathcal{A}}\otimes\tau)(\textup{id}_{\mathcal{A}}\otimes\lambda)\circ\delta$, where $\tau$ is the canonical faithful tracial state on $C_{\lambda}^{*}(G)$. Since $\tau$ is faithful, and $(\textup{id}_{\mathcal{A}}\otimes\lambda)\circ\delta$ is completely isometric by assumption, $\mathbb{E}_{e}$ is faithful.
@@
For the proof of Theorem 3.5, we need a slight strengthening of [DKKLL22, Proposition 3.4] which uses Fell’s absorption principle. See [CD23, § 5.5] and [Ka25] for versions of Fell’s absorption principle for monoid operator algebras.
@@ 2.2. Operator algebras from categories
The operator algebra $\mathcal{A}_{\lambda}(\mathfrak{C})$ always has a contractive self-adjoint approximate identity. Indeed, the net of projections $\sum_{u\in F}\lambda_{u}$ as $F$ ranges over the finite subsets of $\mathfrak{C}^{0}$ is a contractive self-adjoint approximate identity of projections for the dense algebra $\textup{alg}(\{\lambda_{\mathfrak{C}}(c):c\in\mathfrak{C}\})$, so it is also a contractive self-adjoint approximate identity for $\mathcal{A}_{\lambda}(\mathfrak{C})$. Moreover, $\mathcal{A}_{\lambda}(\mathfrak{C})$ is unital whenever $\mathfrak{C}^{0}$ is finite.
@@
where the product $c_{1}^{-1}d_{1}\cdots c_{n}^{-1}d_{n}$ is taken inside $\mathcal{I}(\mathfrak{C})$. We denote by 0 the empty function on $\mathfrak{C}$. Using [CELY17, Lemma 5.6.43], we see that $I_{l}$ contains 0 unless $\mathfrak{C}$ is a monoid such that $x\mathfrak{C}\cap y\mathfrak{C}\neq\emptyset$ for all $x,y\in\mathfrak{C}$. Combining (2.2) and (2.5), we get that
@@
if there exists $X\in\mathcal{J}$ such that $\chi(X)=1$ and $s(x)=t(x)$ for all $x\in X$. We let $s,\chi]$ denote the equivalence class of $(s,\chi)$ with respect to $\sim$. The transformation groupoid $I_{l}\ltimes\Omega$ is the quotient space $(I_{l}*\Omega)/\sim$ with groupoid operations determined by
$s,t.\chi][t,\chi]\coloneq[st,\chi]\quad\text{ and }\quad[s,\chi]^{-1}\coloneq[s^{-1},s.\chi]$
for all $s,t\in I_{l}$ and $\chi\in\Omega$. The range and source maps are given by $\mathfrak{r}([s,\chi])\coloneq s.\chi$ and $\mathfrak{s}([s,\chi])\coloneq\chi$, respectively, for all $s,\chi]\in I_{l}\ltimes\Omega$. Note that Exel calls $I_{l}\ltimes\Omega$ the groupoid of germs for the action $I_{l}\text{\reflectbox{$\uptodownarrow$}}\Omega$, see [Ex08]. We tacitly identify $\Omega$ with the unit space of $I_{l}\ltimes\Omega$ via $\chi\mapsto[\textup{id}_{X},\chi]$ for all $\chi\in\Omega$, where $X\in\mathcal{J}$ is any constructible right ideal with $\chi(X)=1$. Subsets of the form $s,U]\coloneq\{[s,\chi]:\chi\in U\}$ for all $s\in I_{l}$ and compact open subsets $U$ of $\Omega(\textup{dom}(s))$ generate a topology on the groupoid $I_{l}\ltimes\Omega$ which is locally compact, étale (both range and source maps are local homeomorphisms), and ample (there is a basis consisting of compact open bisections) though not necessarily Hausdorff. A bisection is a subset of $I_{l}\ltimes\Omega$ on which both the range and source maps are injective, and any basic open subset $[s,U]$ as above is a compact open bisection.
@@
Let us recall some terminology from [Ex08, Definition 11.5]. A subset $F\subseteq\mathcal{J}$ of constructible right ideals is said to be a cover for $X\in\mathcal{J}$ if $Z\subseteq X$ for all $Z\in F$ and for every $Y\in\mathcal{J}$ with $Y\subseteq X$, there exists $Z\in F$ such that $Z\cap Y\neq\emptyset$. If $0\in I_{l}$, then $\partial\Omega$ is precisely the set of tight characters of $\mathcal{J}$ in the sense of Exel, see [Ex08, Theorem 12.9]. Precisely, this means that $\chi\in\Omega$ lies in $\partial\Omega$ if and only if whenever $X\in\mathcal{J}^{\times}$ and $F$ is a finite cover for $X$, we have $\chi(Y)=1$ for some $Y\in F$. This implies that $\partial\Omega(X,\mathfrak{f})=\emptyset$ whenever $\mathfrak{f}$ is a cover for $X$ (cf. [DM14] and [Ex21]).
@@
for all $f\in C_{c}(I_{l}\ltimes\Omega)$ and $t,\chi]\in(I_{l}\ltimes\Omega)_{\chi}$. Similarly, given $\chi\in\partial\Omega$, we let
@@
determined by $q_{\partial}(f)=f|_{I_{l}\ltimes\partial\Omega}$ for all $f\in C_{c}(I_{l}\ltimes\Omega)$. Note that such a *-homomorphism exists because the restriction map $C_{c}(I_{l}\ltimes\Omega)\to C_{c}(I_{l}\ltimes\partial\Omega)$ is contractive for the reduced C*-norms: the norm on $C_{c}(I_{l}\ltimes\Omega)$ is defined by a supremum over $\Omega$, where as the norm on $C_{c}(I_{l}\ltimes\partial\Omega)$ is defined by a supremum over $\partial\Omega$, a subset of $\Omega$.
When $I_{l}\ltimes\Omega$ is Hausdorff, the isomorphism in (2.7) justifies Definition 2.11. In the non-Hausdorff setting, there is another candidate for a groupoid model for $C_{\lambda}^{*}(\mathfrak{C})$ (see [Li23, § 3]), and these two groupoids can be different, see [Sch+].